OpenStudy (anonymous):

how can you solve cos2x=0

OpenStudy (anonymous):

What angles make cosine =0?

OpenStudy (anonymous):

I actually need a trig recap because I havent taken trig in a year and now i to review it for my ap calc h.w.

OpenStudy (anonymous):

do i use the unit circle?

OpenStudy (anonymous):

I can tell you but it is important you know how cosine function behaves. Google it and come back to me, I am willing to help you!

OpenStudy (anonymous):

You can use the unit circle, of course!

OpenStudy (anonymous):

what do you mean by angles that make up cosine?

OpenStudy (anonymous):

the angle whose cosine function is zero

OpenStudy (anonymous):

well, I have to go. I will give you the solution. Angles that make cosine =0 are:\[\theta=(2k+1)\frac{ \pi }{ 2 }\] where k=[0,±1,±2,±3..........and so on], then your solution is:\[2x=\theta=(2k+1)\frac{ \pi }{ 2 } \rightarrow x=(2k+1)\frac{ \pi }{ 4 }\]

zepdrix (zepdrix):

\[\Large \cos2x=0\]Applying `Cosine Double Angle Formula`:\[\Large 2\cos^2x-1=0\]Solve for cos x! :)

zepdrix (zepdrix):

That's how I would approach it at least :3